OpenStudy (izzy18970):

I need help!!! Write an solve an equation to solve each problem Marty paid $121 for shoes and clothes. He paid $45 more for clothes than he did for shoes. How much did Marty pay for the shoes?

OpenStudy (izzy18970):

can someone pleasee help me!!!!!???????

jimthompson5910 (jim_thompson5910):

let x = amount he paid for shoes only y = amount he paid for clothes only

jimthompson5910 (jim_thompson5910):

we are told that "Marty paid $121 for shoes and clothes" so that must mean that the individual amounts must add to 121, ie x+y = 121

jimthompson5910 (jim_thompson5910):

then we're given that "He paid $45 more for clothes than he did for shoes" which means that (amount paid for clothes) = (amount paid for shoes) + ($45) y = x +45

jimthompson5910 (jim_thompson5910):

now use the two equations to solve for x and y x+y = 121 x+x+45 = 121 .. plug in y = x+45 2x + 45 = 121 x = ???

OpenStudy (izzy18970):

so the equation is 2x + 45 = 21

jimthompson5910 (jim_thompson5910):

it's one of the equations in the steps you use, so yes (in a way)

OpenStudy (izzy18970):

2x + 45 = 21 -45 -45 -24/2 x = -12

OpenStudy (izzy18970):

is this correct?

jimthompson5910 (jim_thompson5910):

oh sorry, it should be 121 and NOT 21

jimthompson5910 (jim_thompson5910):

just caught that error

OpenStudy (izzy18970):

ok let me try again 2x + 45 = 121 -45 -45 76/2 x = 38

jimthompson5910 (jim_thompson5910):

good, then use this to find y

OpenStudy (izzy18970):

38 + 45 = y?

OpenStudy (izzy18970):

is this correct?

jimthompson5910 (jim_thompson5910):

good, keep going

OpenStudy (izzy18970):

83

jimthompson5910 (jim_thompson5910):

so that means x = 38 y = 83

OpenStudy (izzy18970):

so marty paid $38 for the shoes

OpenStudy (izzy18970):

thank you for the help

jimthompson5910 (jim_thompson5910):

you got it and you're welcome